If events A and B are such that they are independent and P(A) = 0.3,
P(B) = 0.5, find
(a) P(An B),
(b) P(A U B).

Answer by ikleyn(52781)

.

(a)  P(A n B) = P(A)*P(B) = 0.3*0.5 = 0.15.    ANSWER


(b)  P(A U B) = P(A) + P(B) - P(A n B) = 0.3 + 0.5 - 0.15 = 0.65.    ANSWER

Solved. // Standard and classic formulas for probabilities.

See any textbook on probability - - - for example, your own textbook.


HINT.   In order for to see formulas in your textbook,  you should  OPEN  your textbook first.

Next,  you should  FIND  appropriate formulas there.